OpenStudy (anonymous):

The point (12, 9) is included in a direct variation. What is the constant of variation? Please explain answer

OpenStudy (amistre64):

direct variation means that y=x and that we need to include a constant "k" to make it even out

OpenStudy (amistre64):

12=9k , solve for k

OpenStudy (amistre64):

err... soo close

OpenStudy (amistre64):

y = xk 9 = 12k thats better :)

OpenStudy (amistre64):

in this case we get: k = 9/12 = 3/4 i believe